The 200 lb box is released from rest at the 1st position and supposed to be at the 2nd position and with no friction on the surface. Theta is 30 degrees.

(a) Determinethevelocity[ft/s]ofthecrateatthe2nd position using the conservation of energy.

(b) Verifytheanswerof(a)usingtheworkandkinetic energy method.

(c) Verify the answer of (a) using the newton
Answer #1

By using the principle of conservation of energy:

mgh=\frac{1}{2}mv^2

v=\sqrt{2gh}

v=\sqrt{2*9.8m/s^2*2m}=1.97m/s

By using the Kinetic work method:

W=\Delta K

mgh=\frac{1}{2}mv^2

v=\sqrt{2gh}

v=\sqrt{2*9.8m/s^2*2m}=1.97m/s

By using the Newton